### v1

# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

Abstract
Underwater visuals undergo various complex degradations, inevitably influencing the efficiency of underwater vision tasks. Recently, diffusion models were employed to underwater image enhancement (UIE) tasks, and gained SOTA performance. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models. In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploiting the knowledge of physics to guide the diffusion process. PA-Diff consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch is a plug-and-play network to produce the physics prior, which can be integrated into any deep framework. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. Extensive experiments prove that our method achieves best performance on UIE tasks.
Index Terms—  Underwater image enhancement, physics model, diffusion model, transformer
1 Introduction
The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics[1] and tracking underwater objects[2]. Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness [3]. Consequently, the acquisition of clear images in underwater scenarios assumes paramount importance, especially in disciplines necessitating interactions with the submerged environment. The primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images. UIE stands as an indispensable component for tasks related to vision in the underwater domain.
To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented [4, 5, 6]. These techniques delve into the physical aspects contributing to degradation, such as color cast or scattering, seeking to rectify and enhance the underwater images. Nevertheless, these models grounded in physics exhibit restricted representational capacities, rendering them insufficient for comprehensively addressing the intricate physical and optical elements inherent in underwater scenes. Consequently, their efficacy diminishes, yielding suboptimal enhancement outcomes, particularly in the face of highly intricate and diverse underwater scenarios. In recent times, many learning-based methodologies [7, 8, 9, 10] have been introduced to yield superior results. Leveraging the potent feature representation and nonlinear mapping capabilities of neural networks, these methods excel. They can effectively learn the transformation from a degraded image to a clear one through extensive paired training data.
Recently, there has been a surge of interest in image synthesis [11, 12] and restoration tasks [13, 14, 15], with a notable focus on diffusion-based techniques like DDPM [16] and DDIM [17]. This heightened attention can be attributed to the remarkable generative capabilities of diffusion models. These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process [16], which poses a new perspective for underwater image enhancement task. Tang et al. [7] present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity. Moreover, the diffusion models lack awareness regarding the difficulty of enhancing different image regions, a critical consideration for modeling the complex distribution in real-world underwater scenes.
In this paper, we develop a novel physics-aware diffusion model, fully exploiting physical information to guide the diffusion process, called PA-Diff. It mainly consists of two branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. The first branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model. The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN) to exploit physics prior information and capture the long-range diffusion dependencies. Our PA-Diff achieves unprecedented perceptual performance in UIE task. Extensive experiments demonstrate that our developed PA-Diff performs the superiority against previous UIE approaches, and ablation study can demonstrate the effectiveness of all contributions.
In summary, the main contributions of our PA-Diff are as follows:
  • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. To the best of our knowledge, it is the first the Diffusion model with the guidance of physical knowledge in underwater image enhancement tasks.
  • We design a physics-aware diffusion transformer block, which not only enables PA-Diff with underwater-aware ability to guide the diffusion process, but also captures the long-range diffusion dependencies.
  • Extensive experiments compared with SOTAs considerably demonstrate that our PA-Diff performs the superiority against previous UIE approaches, and extensive ablation experiments can demonstrate the effectiveness of all contributions.
Fig. 1: Overall framework of our proposed PA-Diff. PA-Diff mainly consists of two cooperative branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch is a plug-and-play universal module to produce the prior knowledge of physics, which can be integrated into any deep learning framework. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
2 Related Works
2.1 Underwater Image Enhancement
Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches [4, 6, 7, 8, 9]. Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors [4], attenuation curve priors [19], fuzzy priors [20]. In addition, Akkaynak and Treibitz [21] proposed a method based on the revised physical imaging model. The manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances. Recently, deep learning-based methods [7, 8, 9] achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance [4, 5, 22]. The previous methods neglect the underwater imaging mechanism and rely only on the representation ability of deep networks. Ucolor [23] combined the underwater physical imaging model in the raw space and designed a medium transmission guided model. ATDCnet [24] attempted to use transmission maps to guide deep neural networks. Therefore, how to fully exploit the knowledge of physics in deep neural networks is a very crucial problem.
2.2 Diffusion Model
Diffusion Probabilistic Models (DPMs) [16, 17] have been widely adopted for conditional image generation [13, 14, 25, 26, 27]. Saharia et al. [28] proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration. Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration [29] and low light image enhancement [30]. Recently, Tang et al. [7] presented an image enhancement approach with diffusion model in underwater scenes. WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
3 Methodology
3.1 Overall Framework
The overall framework of PA-Diff is shown in Fig. 1. PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model. The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. We adopt the diffusion process proposed in DDPM [16] to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
Forward Diffusion Process. The forward diffusion process can be viewed as a Markov chain progressively adding Gaussian noise to the data. Given a clean data $x_{0}$, then introduce Gaussian noise based on the time step, as follows:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (1)
where $\beta_{t}$ is a variable controlling the variance of the noise. Introducing $\alpha_{t}=1-\beta_{t}$, this process can be described as:
$(x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).\text{(}$2)
With Gaussian distributions are merged, We can obtain :
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$ (3)
Reverse Diffusion Process. The reverse diffusion process aims to restore the clean data from the Gaussian noise. The reverse diffusion can be expressed as:
$p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$ (4)
where $x_{c}$ refers to the conditional image ${I}$ (underwater input image). $\boldsymbol{\mu}_{\theta}(x_{t},x_{c},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively. We follow the setup of [16], they can be expressed as:
$\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$ (5)
$\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$ (6)
where $\epsilon_{\theta}(x_{t},x_{c},t)$ is the estimated value with a Unet.
We optimize an objective function for the noise estimated by the network and the noise $\epsilon$ actually added. Therefore, the diffusion loss is:
$L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$ (7)
3.2 Physics Prior Generation
Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model:
$I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$ (8)
where $c\in\{R,G,B\}$ is the color channel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}(x)$ is the medium transmission map, and $B^{c}$ is the global background scattered light. In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}(x)$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
Network architecture of PPG branch is shown in Fig. 1, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$. The two sub-networks consist of duplicated Dynamic Convolutions (DC) [31]. DC dynamically combines numerous parallel convolutional kernels according to their attention, thereby enhancing the model’s complexity and performance without augmenting the network’s depth or breadth.
Given a underwater image as input $I$ and its corresponding ground truth (GT), we can obtain via Eq. 8:
$\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(I),$ (9)
We constrain the PPG branch by supervised manner. The physics reconstruction loss can be expressed as:
$\mathcal{L}_{p}=||\hat{I}-I||_{1},$ (10)
Finally, we use a Conv and GAP to extract physics prior information $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ to guide the diffusion process.
3.3 Physics-aware Diffusion Transformer
Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration. Fig. 1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer blocks (PDTB). We also utilize skip connections to connect the features at the same level.
Given a degraded underwater image as condition image $x_{c}\in{R}^{H\times W\times 3}$, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch. Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel. $\mathbf{F}$ is encoded and decoded by PDTB, which consists of physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN). In our PA-MSA, We first embed the time embedding $\mathbf{T}$ into the input features $\mathbf{F}$ :
$\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{T},$ (11)
where $Norm$ means layer normalization. Then, we integrated $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ as the corresponding dynamic modulation parameters to generate transmission-aware feature $\mathbf{F_{t}}$ and light-aware feature $\mathbf{F_{b}}$:
$\mathbf{F_{t}}$ $=Li(\mathbf{T^{c}})\odot\tilde{\mathbf{F}},$ (12)
$\mathbf{F_{b}}$ $=Li(\mathbf{B^{c}})\odot\tilde{\mathbf{F}},$
where $Li()$ means linear layer. Afterward, we aggregate the obtained embeddings by projecting $\mathbf{F_{b}}$ into query $\mathbf{Q}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{b}}$, $\mathbf{F_{t}}$ into key $\mathbf{K}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{t}}$ and transforming $\tilde{\mathbf{F}}$ into value $\mathbf{V}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution. We employ self-attention and get the output of PA-SA $\mathbf{\hat{F}}$:
$\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$ (13)
where $\alpha$ is a learnable parameter. Consequently, PA-SA introduces physical guidance to fully exploit physics knowledge at the feature level and use self-attention mechanism to implicitly model the features of transmission map and background scattered light, which can help the diffusion model restore missing details and correct color distortion.
Finaly, we design a multi-scale dynamic feed-forward network (MS-FFN) for local feature aggregation. In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions. MS-FFN adopts GELU to ensure the flexibility of feature aggregation. Thus, the multi-scale feature $\mathbf{F^{3}}$ and $\mathbf{F^{5}}$ of MS-FFN can be expressed as:
$\mathbf{F^{3}}$ $=\mathbf{W}^{3}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{3}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$ (14)
$\mathbf{F^{5}}$ $=\mathbf{W}^{5}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{5}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$
The output feature $F_{out}$ of MS-FFN can be expressed as:
$\mathbf{F_{out}}=Conv(Concat(\mathbf{F^{3}},\mathbf{F^{5}}))+\hat{\mathbf{F}}.$ (15)
Table 1: Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
Methods UIEWD UWCNN UIECˆ2-Net Water-Net U-color U-shape DM-water Ours
[…]
Fig. 2: Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
4 Experiments
4.1 Experimental Settings
Implementation details. Our network, implemented using PyTorch 1.7, underwent training and testing on an NVIDIA GeForce RTX 3090 GPU. We employed the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was established at 0.0001. During the training stage, to balance the batch size and image size, the batch and image size are set to 8 and 128 × 128, respectively. The number of training iterations reached one million. The pixel values of the image are normalized to [-1,1]. The time step of the diffusion model is set to 2000. In the testing stage, following the setting in [7], the input size of the image is 256 × 256. By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
Datasets. We utilize the real-world UIEBD dataset [8] and the LSUI dataset [9] for training and evaluating our model. The UIEBD dataset comprises 890 underwater images with corresponding labels. Out of these, 700 images are allocated for training, and the remaining 190 are designated for testing. LSUI dataset contains 5004 underwater images and their corresponding high-quality images. Compared with UIEBD, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images. In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing.
Evaluation Metrics. We employ standard full-reference image quality assessment metrics, including PSNR and SSIM [32], for quantitative comparisons at both pixel and structural levels. Higher PSNR and SSIM values indicate superior image quality. Additionally, LPIPS [33] and FID [34] are utilized to evaluate perceptual performance. Lower LPIPS and FID scores signify a more effective UIE approach.
4.2 Results and Comparisons
Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD and LSUI datasets, including with UIECˆ2-Net [35], Water-Net [8], UIEWD [36], UWCNN [37], U-color [23], U-shape [9], and DM-water [7]. We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets. The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics. To better validate the superiority of our methods, Fig. 2 shows the visual comparison results with other methods. The six examples are randomly selected on the UIEBD and LSUI datasets. Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
4.3 Ablation Study
To evaluate the impact of each strategy on the diffusion model, we conduct an ablation study. Table 2 shows the ablation results on the LSUI dataset. Model A means that we just use the normal transformer structure. Compared with model A, model B, C and D achieve better results, suggesting that the knowledge of physics can help diffusion model to better restore image. Model B outperforms C, suggesting that the prior information of the transmission map is more important than the background light. Model E achieves the best performance, proving that our all contribution is effective for UIE tasks. Model E underperforms model F. The possible reason for this is that there is still a gap between the generated physics information (the transmission map and the background light) and the ideal true value. Therefore, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
Table 2: Ablation study on LSUI dataset. $T$ and $B$ refer to transmission map and background light prior information, respectively. $MS$ means multi-scale operation in our designed MS-FFN. $GAP$ denotes that we use global average pooling for the extraction of physical prior information from both the generated transmission map and background light.
[…]
A $\times$ $\times$ $\times$ $\times$ 28.35 0.8943
B $\checkmark$ $\times$ $\times$ $\checkmark$ 24.71 0.9193
C $\times$ $\checkmark$ $\times$ $\checkmark$ 25.84 0.9077
D $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 23.06 0.9298
E $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 25.44 0.9177
F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 22.15 0.9354
5 Conclusion
In this paper, we develop a novel UIE framework, namely PA-Diff. With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. To the best of our knowledge, it is the first the Diffusion model based on physical perception in image enhancement tasks. Our designed physics prior generation branch is a plug-and-play universal module to produce physics prior information, which can be integrated into any deep learning framework. PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.

### v2

@@
 Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement
-Abstract
+Abstract.
 Underwater visuals undergo various complex degradations, inevitably influencing the efficiency of underwater vision tasks.
@@
 In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploiting the knowledge of physics to guide the diffusion process.
-PA-Diff consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch.
-Our designed PPG branch is a plug-and-play network to produce the physics prior, which can be integrated into any deep framework.
+PA-Diff consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
+Our designed PPG branch aims to produce the prior knowledge of physics.
 With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
+INR Branch can learn robust feature representations from diverse underwater image via implicit neural representation, which reduces the difficulty of restoration for PDT branch.
 Extensive experiments prove that our method achieves best performance on UIE tasks.
-Index Terms—  Underwater image enhancement, physics model, diffusion model, transformer
-1 Introduction
-The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics[1] and tracking underwater objects[2].
-Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness [3].
-Consequently, the acquisition of clear images in underwater scenarios assumes paramount importance, especially in disciplines necessitating interactions with the submerged environment.
-The primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images.
-UIE stands as an indispensable component for tasks related to vision in the underwater domain.
-To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented [4, 5, 6].
+The code is available at https://github.com/chenydong/PA-Diff.
+Underwater image enhancement, Physics model, Diffusion model, Transformer
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer vision
+1. Introduction
+The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics(McMahon and Plaku, 2021) and tracking underwater objects(de Langis and Sattar, 2020).
+Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness (Akkaynak et al., 2017).
+Consequently, the primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images.
+To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented (Peng and Cosman, 2017; Jr.
+et al., 2013; Peng et al., 2018).
 These techniques delve into the physical aspects contributing to degradation, such as color cast or scattering, seeking to rectify and enhance the underwater images.
@@
 Consequently, their efficacy diminishes, yielding suboptimal enhancement outcomes, particularly in the face of highly intricate and diverse underwater scenarios.
-In recent times, many learning-based methodologies [7, 8, 9, 10] have been introduced to yield superior results.
+In recent times, many learning-based methodologies (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fabbri et al., 2018) have been introduced to yield superior results.
 Leveraging the potent feature representation and nonlinear mapping capabilities of neural networks, these methods excel.
-They can effectively learn the transformation from a degraded image to a clear one through extensive paired training data.
-Recently, there has been a surge of interest in image synthesis [11, 12] and restoration tasks [13, 14, 15], with a notable focus on diffusion-based techniques like DDPM [16] and DDIM [17].
-This heightened attention can be attributed to the remarkable generative capabilities of diffusion models.
-These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process [16], which poses a new perspective for underwater image enhancement task.
+Deep learning-based methods can be roughly categorized into two classes: physics-free end-to-end methods (Jiang et al., 2022; Peng et al., 2023) and physics-aware methods (Li et al., 2021; Mu et al., 2023b; Fu et al., 2022a).
+Regarding the physics-free end-to-end methods, most of them usually use ground-truth images as labels to enforce L1/L2 distance-based consistency and also introduce various regularizations as additional constraints to deal with the ill-posed property.
+Although these methods achieve decent performance, they mainly consider the model in the raw image space, which does not fully explore the useful physics-aware feature information for UIE.
+Regarding the physics-aware deep methods, most of them utilize the atmospheric scattering model.
+Existing UIE approaches (Li et al., 2021; Fu et al., 2022a) develop deep neural networks to estimate the transmission and the background scattered light, and then follow the traditional UIE methods to restore underwater images.
+However, these approaches do not correct the errors of the atmospheric light estimated by neural networks, which seriously affects the quality of the images (Ye et al., 2022).
+Therefore, there is a crucial problem to adequately exploit the physical information and effectively incorporate them into a unified UIE deep framework.
+Recently, there has been a surge of interest in image synthesis (Rombach et al., 2022; Saharia et al., 2022b; Lu et al., 2024, 2023; Zhou et al., 2024) and restoration tasks (Choi et al., 2021; Wang et al., 2023; Yi et al., 2023; Zhou et al., 2023), with a notable focus on diffusion-based techniques like DDPM (Ho et al., 2020) and DDIM (Song et al., 2021).
+This heightened attention can be attributed to the remarkable generative capabilities of diffusion models.These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process (Ho et al., 2020), which poses a new perspective for underwater image enhancement task.
 Tang et al.
-[7] present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
-However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity.
+(Tang et al., 2023) present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
+UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023).
+However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity for the diffusion model.
 Moreover, the diffusion models lack awareness regarding the difficulty of enhancing different image regions, a critical consideration for modeling the complex distribution in real-world underwater scenes.
 In this paper, we develop a novel physics-aware diffusion model, fully exploiting physical information to guide the diffusion process, called PA-Diff.
-It mainly consists of two branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch.
-The first branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model.
+It mainly consists of three branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
+The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model (Akkaynak and Treibitz, 2019a).
 The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability.
+The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch.
+Inspired by the latest works (Chen et al., 2024; Yang et al., 2023a) based on INR that can encode an image as a continuous function and learn a more robust degradation representation, we further incorporate the INR into our proposed framework to learn robust feature representations from diverse underwater image.
 The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information.
-Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN) to exploit physics prior information and capture the long-range diffusion dependencies.
+Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA), physics perception unit (PPU) and gated multi-scale feed-forward network (GM-FFN) to exploit physics prior information and capture the long-range diffusion dependencies.
 Our PA-Diff achieves unprecedented perceptual performance in UIE task.
@@
 In summary, the main contributions of our PA-Diff are as follows:
-  • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch.
+  • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
 To the best of our knowledge, it is the first the Diffusion model with the guidance of physical knowledge in underwater image enhancement tasks.
+  • We incorporate Implicit Neural Reconstruction into our proposed PA-Diff to learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
   • We design a physics-aware diffusion transformer block, which not only enables PA-Diff with underwater-aware ability to guide the diffusion process, but also captures the long-range diffusion dependencies.
+Moreover, we also propose a physics perception unit (PPU), which can explore the beneficial feature-level physics information for underwater image enhancement.
   • Extensive experiments compared with SOTAs considerably demonstrate that our PA-Diff performs the superiority against previous UIE approaches, and extensive ablation experiments can demonstrate the effectiveness of all contributions.
-Fig.
-1: Overall framework of our proposed PA-Diff.
-PA-Diff mainly consists of two cooperative branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch.
-Our designed PPG branch is a plug-and-play universal module to produce the prior knowledge of physics, which can be integrated into any deep learning framework.
+Figure 1.
+Overall framework of our proposed PA-Diff.
+PA-Diff mainly consists of three cooperative branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
+Our designed PPG branch aims to produce the prior knowledge of physics.
 With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
-2 Related Works
-2.1 Underwater Image Enhancement
-Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches [4, 6, 7, 8, 9].
-Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors [4], attenuation curve priors [19], fuzzy priors [20].
-In addition, Akkaynak and Treibitz [21] proposed a method based on the revised physical imaging model.
-The manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances.
-Recently, deep learning-based methods [7, 8, 9] achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance [4, 5, 22].
+INR Branch can learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
+2. Related Works
+2.1. Underwater Image Enhancement
+Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches (Peng and Cosman, 2017; Peng et al., 2018; Tang et al., 2023; Peng et al., 2023; Li et al., 2020).
+Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors (Peng and Cosman, 2017), attenuation curve priors (Wang et al., 2018), fuzzy priors (Chiang and Chen, 2012).
+In addition, Akkaynak and Treibitz (Akkaynak and Treibitz, 2019b) proposed a method based on the revised physical imaging model.
+However, the depth map of the underwater scene is difficult to obtain.
+This leads to unstable performance of the this method, which usually suffers from severe color cast and artifacts.
+Therefore, the manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances.
+In recent years, underwater physical imaging models are gradually utilized in combination with data-driven methods (Li et al., 2021).
+Recently, deep learning-based methods (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fu et al., 2022b) achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance (Peng and Cosman, 2017; Jr.
+et al., 2013; Zhao et al., 2023).
 The previous methods neglect the underwater imaging mechanism and rely only on the representation ability of deep networks.
-Ucolor [23] combined the underwater physical imaging model in the raw space and designed a medium transmission guided model.
-ATDCnet [24] attempted to use transmission maps to guide deep neural networks.
+Ucolor (Li et al., 2021) combined the underwater physical imaging model in the raw space and designed a medium transmission guided model.
+ATDCnet (Mu et al., 2023a) attempted to use transmission maps to guide deep neural networks.
 Therefore, how to fully exploit the knowledge of physics in deep neural networks is a very crucial problem.
-2.2 Diffusion Model
-Diffusion Probabilistic Models (DPMs) [16, 17] have been widely adopted for conditional image generation [13, 14, 25, 26, 27].
+2.2. Diffusion Model
+Diffusion Probabilistic Models (DPMs) (Ho et al., 2020; Song et al., 2021) have been widely adopted for conditional image generation (Choi et al., 2021; Wang et al., 2023; Lu et al., 2023; Zhou et al., 2023, 2024).
 Saharia et al.
-[28] proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration.
-Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration [29] and low light image enhancement [30].
+(Saharia et al., 2022a) proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration.
+Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration (Özdenizci and Legenstein, 2023), document enhancement(Yang et al., 2023b), and low light image enhancement (Yang et al., 2023c; Zhou et al., 2023).
 Recently, Tang et al.
-[7] presented an image enhancement approach with diffusion model in underwater scenes.
-WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
+(Tang et al., 2023) presented an image enhancement approach with diffusion model in underwater scenes.
+WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
+UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023).
 However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
-3 Methodology
-3.1 Overall Framework
+2.3. Implicit Neural Representation
+Implicit Neural Representations (INRs) have emerged as an innovative technique garnering substantial attention for their ability to utilize coordinate-based Multi-Layer Perceptrons (MLPs) to model continuous-domain signals.
+Recently, some works have delved into the potential of INRs within the realm of two-dimensional image processing, including applications like image compression (Strümpler et al., 2022), and super-resolution(Chen et al., 2023).
+Chen et al.
+(Chen et al., 2021) first capitalized on INRs to achieve continuous image super-resolution.
+Subsequently, Lee et al.
+(Lee et al., 2022) proposed a dominant-frequency estimator that effectively predicts local texture information in natural images.
+Sun et al.
+(Sun et al., 2021) leveraged INRs to forecast continuous attributes based on the captured tomographic features.
+More recently, Quan et al.
+(Quan et al., 2023) introduced an INR-based inverse kernel prediction network to address the problem of single image defocus deblurring.
+Concurrently, Yang et al.
+(Yang et al., 2023a) exploited the controllable fitting capabilities of INRs to tackle the challenge of low-light image enhancement.
+Chen et al.
+(Chen et al., 2024) utilized INRs to learn common rain degradation representations from diverse rainy scenes.
+These efforts (Sun et al., 2021; Yang et al., 2023a; Chen et al., 2024) convincingly demonstrate the robust representation capabilities of INRs when applied to degraded images.
+Thus, inspired by the previous works, we explore the application of INRs to obtain robust feature representations for underwater image enhancement, aiming to reduce the difficulty of restoration for diffusion models.
+Figure 2. The detailed architecture of our designed (a) physics-aware self-attention (PA-SA) and (b) gated multi-scale feed-forward network (GM-FFN).
+3. Methodology
+3.1. Overall Framework
 The overall framework of PA-Diff is shown in Fig.
 1.
-PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch.
+PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
 The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model.
 The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability.
+The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch.
+Inspired by (Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce implicit neural representations to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models.
 The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information.
-We adopt the diffusion process proposed in DDPM [16] to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
+We adopt the diffusion process proposed in DDPM (Ho et al., 2020) to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
 Forward Diffusion Process.
@@
 Given a clean data $x_{0}$, then introduce Gaussian noise based on the time step, as follows:
-$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (1)
+(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
 where $\beta_{t}$ is a variable controlling the variance of the noise. Introducing $\alpha_{t}=1-\beta_{t}$, this process can be described as:
-$(x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).\text{(}$2)
+(2) $x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).$
 With Gaussian distributions are merged, We can obtain :
-$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$ (3)
+(3) $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$
 Reverse Diffusion Process. The reverse diffusion process aims to restore the clean data from the Gaussian noise. The reverse diffusion can be expressed as:
-$p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$ (4)
+(4) $p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$
 where $x_{c}$ refers to the conditional image ${I}$ (underwater input image).
 $\boldsymbol{\mu}_{\theta}(x_{t},x_{c},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively.
-We follow the setup of [16], they can be expressed as:
-$\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$ (5)
-$\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$ (6)
+We follow the setup of (Ho et al., 2020), they can be expressed as:
+(5) $\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$
+(6) $\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$
 where $\epsilon_{\theta}(x_{t},x_{c},t)$ is the estimated value with a Unet.
 We optimize an objective function for the noise estimated by the network and the noise $\epsilon$ actually added. Therefore, the diffusion loss is:
-$L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$ (7)
-3.2 Physics Prior Generation
-Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model:
-$I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$ (8)
-where $c\in\{R,G,B\}$ is the color channel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}(x)$ is the medium transmission map, and $B^{c}$ is the global background scattered light.
-In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}(x)$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
+(7) $L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$
+3.2. Physics Prior Generation
+Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model (Akkaynak and Treibitz, 2019a):
+(8) $I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$
+where $c\in\{R,G,B\}$ is the color channel, $x$ indicates the spatial location of each pixel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}$ is the medium transmission map, and $B^{c}$ is the global background scattered light.
+In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
 Network architecture of PPG branch is shown in Fig.
 1, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$.
-The two sub-networks consist of duplicated Dynamic Convolutions (DC) [31].
+The two sub-networks consist of duplicated Dynamic Convolutions (DC) (Chen et al., 2020).
 DC dynamically combines numerous parallel convolutional kernels according to their attention, thereby enhancing the model’s complexity and performance without augmenting the network’s depth or breadth.
+Given that the background light differs fundamentally from the transmission map, being independent of image content and instead encapsulating the global properties (Fu et al., 2022a) , for the purpose of generating the background light component, our background light generation sub-network initially employs a Gaussian blur block (GBB) to eliminate fine-grained content details and thereby extract the inherent global attributes.
 Given a underwater image as input $I$ and its corresponding ground truth (GT), we can obtain via Eq. 8:
-$\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(I),$ (9)
-We constrain the PPG branch by supervised manner. The physics reconstruction loss can be expressed as:
-$\mathcal{L}_{p}=||\hat{I}-I||_{1},$ (10)
-Finally, we use a Conv and GAP to extract physics prior information $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ to guide the diffusion process.
-3.3 Physics-aware Diffusion Transformer
+(9) $\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(GBB(I)),$
+We constrain the PPG branch by supervised manner. The physics reconstruction loss of PPG branch can be expressed as:
+(10) $\mathcal{L}_{p}=||\hat{I}-I||_{1}.$
+To enhance the perceptual similarity of the reconstructed images, we introduce the perceptual Loss $\mathcal{L}_{per}$ :
+(11) $\mathcal{L}_{per}=||\phi(\hat{I})-\phi(I)||_{2},$
+where $\phi$ is a pre-trained VGG network.
+Finally, the overall loss $\mathcal{L}_{ppg}$ of PPG branch can be expressed as:
+(12) $\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$
+Figure 3. The detailed structure of the proposed physics perception unit (PPU).
+3.3. Implicit Neural Reconstruction
+Inspired by (Quan et al., 2023; Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models.
+INRs trains a multi-layer perceptron (MLP) by learning the following mapping :
+(13) $F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$
+where the dimensions of input refer to the (x, y) spatial coordinates, while the dimensions of output refer to the (R, G, B) color channels.
+Specifically, following the setting of (Chen et al., 2024; Yang et al., 2023a), underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with a image resolution of H × W.
+Meanwhile, the position of each pixel is recorded in a relative coordinate set $\mathbf{X}\in\mathbb{R}^{H\times W\times 2}$, where the value ’2’ means horizontal and vertical coordinates.
+As shown in Figure 1, we fuse $X$ and $E$, and use $F_{\theta}$ to output image $\boldsymbol{I}_{{\mathrm{inr}}}$, which is expressed as:
+(14) $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$
+where $[x,y]$ is the location of a pixel and $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]$ is the generated RGB value.
+Due to the elimination of high-frequency components during rendering (Yang et al., 2023a), we adopt periodic spatial encoding (Lee and Jin, 2022) to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details.
+The encoding procedure is formulated as:
+(15) $\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$
+(16) $\gamma(c)=(\cdot\cdot\cdot,\sin(2^{i}c),\cos(2^{i}c),\cdot\cdot\cdot),$
+where $\gamma()$ represents a spatial encoding function.
+$i$ values from 00 to $L-1$, $c\in\mathbb{R}^{2L}$ is the coordinate value of $X$, and $L$ is a hyperparameter that determines the dimension value.
+Given a GT of underwater image $GT$, The reconstruction loss of INR branch can be expressed as:
+(17) $\mathcal{L}_{INR}=||\boldsymbol{I}_{\mathrm{inr}}-GT||_{1}.$
+Finally, we fuse $I$ and $\boldsymbol{I}_{\mathrm{inr}}$ to obtain the condition $x_{c}$ of diffusion model, which can be expressed as:
+(18) $x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$
+3.4. Physics-aware Diffusion Transformer
 Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration.
 Fig.
-1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer blocks (PDTB).
-We also utilize skip connections to connect the features at the same level.
-Given a degraded underwater image as condition image $x_{c}\in{R}^{H\times W\times 3}$, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch.
+1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention module (CAM).
+Given the output of the INR branch as condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch.
 Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel.
-$\mathbf{F}$ is encoded and decoded by PDTB, which consists of physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN).
-In our PA-MSA, We first embed the time embedding $\mathbf{T}$ into the input features $\mathbf{F}$ :
-$\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{T},$ (11)
+$\mathbf{F}$ is encoded and decoded by PDT block, which consists of physics-aware self-attention (PA-SA), physics perception unit and gated multi-scale feed-forward network (GM-FFN).
+Note that, we introduce a physics perception unit (PPU) to further exploit the beneficial feature-level physics information.
+Cross Attention Module.
+CAM aims to obtain a physics-aware interaction embedding, using which guides the diffusion process.
+We denote $T^{c}$ and $\mathbf{F}$ as the input of CAM.
+We use different linear projections to construct $Q$ and $K$ in CAM:
+(19) $Q=Conv_{1\times 1}(T^{c}),$
+(20) $K=Conv_{1\times 1}(\mathbf{F}).$
+Similarly, $V$ of the transmission map $T^{c}$ can be obtained:
+(21) $V=Conv_{1\times 1}(T^{c}).$
+The output feature map of CAM $T_{out}$ can then be obtained from the formula:
+(22) $T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$
+Physics-aware Self-attention. As is shown in Figure 2 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA:
+(23) $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t},$
 where $Norm$ means layer normalization.
-Then, we integrated $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ as the corresponding dynamic modulation parameters to generate transmission-aware feature $\mathbf{F_{t}}$ and light-aware feature $\mathbf{F_{b}}$:
-$\mathbf{F_{t}}$ $=Li(\mathbf{T^{c}})\odot\tilde{\mathbf{F}},$ (12)
-$\mathbf{F_{b}}$ $=Li(\mathbf{B^{c}})\odot\tilde{\mathbf{F}},$
-where $Li()$ means linear layer.
-Afterward, we aggregate the obtained embeddings by projecting $\mathbf{F_{b}}$ into query $\mathbf{Q}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{b}}$, $\mathbf{F_{t}}$ into key $\mathbf{K}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{t}}$ and transforming $\tilde{\mathbf{F}}$ into value $\mathbf{V}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution.
+We first obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$.
+Meanwhile, we generate query ($Q_{t}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{T_{out}}$) from the output feature map of CAM, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution.
+We employ physics-aware query-driven attention strategy to obtain the final query $Q$ by fusing $Q_{t}$ and $Q_{f}$, which can be expressed as:
+(24) $Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$
 We employ self-attention and get the output of PA-SA $\mathbf{\hat{F}}$:
-$\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$ (13)
+(25) $\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$
 where $\alpha$ is a learnable parameter.
-Consequently, PA-SA introduces physical guidance to fully exploit physics knowledge at the feature level and use self-attention mechanism to implicitly model the features of transmission map and background scattered light, which can help the diffusion model restore missing details and correct color distortion.
-Finaly, we design a multi-scale dynamic feed-forward network (MS-FFN) for local feature aggregation.
+Consequently, PA-SA introduces physical guidance to fully exploit physics knowledge at the feature level and use self-attention mechanism to implicitly model the features of transmission map, which can help the diffusion model restore missing details and correct color distortion.
+Physics Perception Unit.
+Exploring feature-level physics information (Zheng et al., 2023; Dong and Pan, 2020) achieved success in image dehazing tasks.
+Our proposed physics perception unit (PPU) aims to further exploit the beneficial feature-level physics information for underwater image enhancement, and learn richer physical features.
+Our proposed PPU is based on the simplified atmospheric model Eq.
+(8), as shown in Fig.
+3.
+We can reformulate the physics model to represent the clear image J, which can be obtained by:
+(26) $J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$
+Let $k$ denote a feature extractor(the filter kernel in a deep network). Then extracting features via the feature extractor k, Eq. (25) can be reformulated as follows:
+(27) $k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$
+where $\otimes$ denotes the convolution operator and $\odot$ denotes the element-wise product operation. Using the matrix-vector form for algebraic operations, Eq. (26) can be express as:
+(28) $\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$
+where $\boldsymbol{K},\boldsymbol{J},\boldsymbol{I},\boldsymbol{B},\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ respectively denote the matrix-vector forms of $k,J,I,B,\frac{1}{t}$ and $(1-\frac{1}{t})$.
+Note that the diagonal vector of the diagonal matrix $\boldsymbol{T}_{1}$ corresponds to the vectorized form of $\frac{1}{t}$.
+Assuming that $\boldsymbol{K}$ is an invertible matrix-vector, we can obtain
+(29) $\boldsymbol{F}_{1}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1},$
+(30) $\boldsymbol{F}_{2}=\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1}.$
+And then, Eq. (27) can be express as:
+(31) $\boldsymbol{KJ}$ $=(\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{I}+(\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{B}$
+$=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$
+The Eq.
+(30) presents the relation between the clear image and the underwater image in a deep feature space.
+Due to the strong representation ability of deep neural networks, we can design a feature-level physics perception unit (PPU) and adopt deep networks to approximate the features corresponding to $\boldsymbol{F}_{1}$, $\boldsymbol{F}_{2}$ and $\boldsymbol{K}\boldsymbol{B}$.
+We denote $\tilde{\boldsymbol{B}}$ as an approximation of the features $\boldsymbol{K}\boldsymbol{B}$ that correspond to the background light, and $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ as approximations of $\boldsymbol{F}_{1}$ and $\boldsymbol{F}_{2}$, which are associated with the transmission map.
+Furthermore, $\boldsymbol{K}\boldsymbol{I}$ and $\boldsymbol{KJ}$ can be viewed as a latent underwater image $\tilde{\boldsymbol{I}}$ and its corresponding latent clear image $\tilde{\boldsymbol{J}}$, respectively.
+According to Eq.
+(30), the $\tilde{\boldsymbol{J}}$ can be express as:
+(32) $\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}.$
+Since the background light is usually assumed to be homogeneous, we can use global average pooling (GAP) to obtain the main information for $\tilde{\boldsymbol{B}}$ in the feature space.
+$\tilde{\boldsymbol{B}}$ can be express:
+(33) $\tilde{\boldsymbol{B}}=H(\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{GAP}(\tilde{\boldsymbol{I}}))))))))),$
+where $\sigma$ is the Sigmoid function, H denotes a replication operation.
+As the transmission map is non-homogeneous, we cannot apply GAP for the approximations of $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$.
+We use different projection networks to obtain $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ :
+(34) $\tilde{\boldsymbol{t}_{1}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))),$
+(35) $\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))).$
+where s represents the shared Conv layer. Given $\hat{\mathbf{F}}$ as the input of PPU, we use residual connection to obtain the final output of PPU, which can be expressed as :
+(36) $\hat{\mathbf{F_{ppu}}}=PPU(\hat{\mathbf{F}})+\hat{\mathbf{F}}.$
+Gated Multi-scale Feed-Forward Network.
+Finaly, we design a GM-FFN for local feature aggregation.
+As is shown in Figure 2 (b), we employ GD-FFN-like structure (Zamir et al., 2022) to construct our feed-forward network.
 In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions.
-MS-FFN adopts GELU to ensure the flexibility of feature aggregation.
-Thus, the multi-scale feature $\mathbf{F^{3}}$ and $\mathbf{F^{5}}$ of MS-FFN can be expressed as:
-$\mathbf{F^{3}}$ $=\mathbf{W}^{3}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{3}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$ (14)
-$\mathbf{F^{5}}$ $=\mathbf{W}^{5}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{5}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$
-The output feature $F_{out}$ of MS-FFN can be expressed as:
-$\mathbf{F_{out}}=Conv(Concat(\mathbf{F^{3}},\mathbf{F^{5}}))+\hat{\mathbf{F}}.$ (15)
-Table 1: Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
+GM-FFN adopts GELU to ensure the flexibility of feature aggregation.
+Figure 4. The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
+Table 1. Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
 Methods UIEWD UWCNN UIECˆ2-Net Water-Net U-color U-shape DM-water Ours
@@
 […]
-Fig. 2: Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
-4 Experiments
-4.1 Experimental Settings
+[…]
+[…]
+Figure 5. Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
+Figure 6. More visual results from real-world scenarios.
+Figure 7. Limitations. The output images refer to the reconstruction images of PPG branch.
+4. Experiments
+4.1. Experimental Settings
 Implementation details.
@@
 The learning rate was established at 0.0001.
-During the training stage, to balance the batch size and image size, the batch and image size are set to 8 and 128 × 128, respectively.
+During the training stage, to balance the batch size and image size, the batch and image size are set to 2 and 256 × 256, respectively.
 The number of training iterations reached one million.
@@
 The time step of the diffusion model is set to 2000.
-In the testing stage, following the setting in [7], the input size of the image is 256 × 256.
+In the testing stage, following the setting in (Tang et al., 2023), the input size of the image is 256 × 256.
 By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
 Datasets.
-We utilize the real-world UIEBD dataset [8] and the LSUI dataset [9] for training and evaluating our model.
+We utilize the real-world UIEBD dataset (Peng et al., 2023) and the LSUI dataset (Li et al., 2020) for training and evaluating our model.
 The UIEBD dataset comprises 890 underwater images with corresponding labels.
@@
 In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing.
+In addition, to verify the generalization of PA-Diff, we use non-reference benchmarks U45 (Li et al., 2019), which contains 45 underwater images for testing.
+Comparison methods.
+We conduct a comparative analysis between PA-Diff and seven state-of-the-art (SOTA) UIE methods, namely UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UWCNN (Anwar et al., 2018), UIEWD (Ma and Oh, 2022), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023).
+To ensure a fair and rigorous comparison, we utilize the provided source codes from the respective authors and adhere strictly to the identical experimental settings across all evaluations.
 Evaluation Metrics.
-We employ standard full-reference image quality assessment metrics, including PSNR and SSIM [32], for quantitative comparisons at both pixel and structural levels.
+We employ standard full-reference image quality assessment metrics, including PSNR and SSIM (Wang et al., 2004), for quantitative comparisons at both pixel and structural levels.
 Higher PSNR and SSIM values indicate superior image quality.
-Additionally, LPIPS [33] and FID [34] are utilized to evaluate perceptual performance.
+Additionally, LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) are utilized to evaluate perceptual performance.
 Lower LPIPS and FID scores signify a more effective UIE approach.
-4.2 Results and Comparisons
-Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD and LSUI datasets, including with UIECˆ2-Net [35], Water-Net [8], UIEWD [36], UWCNN [37], U-color [23], U-shape [9], and DM-water [7].
-We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets.
+For non-reference benchmarks U45, we introduce UIQM (Panetta et al., 2015) and UCIQE (Yang and Sowmya, 2015) to evaluate our method.
+Table 2.
+Ablation study on LSUI dataset.
+$T^{c}$ and $B^{c}$ refer to transmission map and background light prior information in PPG branch, respectively.
+GBB means Gaussian blur module.
+INR represents our INR branch.
+PSE refers to the periodic spatial encoding (Lee and Jin, 2022), aiming to better recover high-frequency details.
+CAM and PPU means our designed cross attention module and physics perception unit , respectively.
+$GM-FFN$ means multi-scale operation in the feed-forward network.
+[…]
+A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 28.35 0.8893
+B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 27.71 0.8974
+C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.92 0.9071
+D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.24 0.9112
+E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 24.50 0.9186
+F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 24.92 0.9176
+G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 23.66 0.9241
+H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 22.85 0.9374
+Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 22.13 0.9378
+4.2. Results and Comparisons
+Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD, LSUI and U45 datasets, including with UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UIEWD (Ma and Oh, 2022), UWCNN (Anwar et al., 2018), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023).
+We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets, and UIQM and UCIQE for non-reference dataset U45.
 The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics.
 To better validate the superiority of our methods, Fig.
-2 shows the visual comparison results with other methods.
-The six examples are randomly selected on the UIEBD and LSUI datasets.
+5 shows the visual comparison results with other methods on the UIEBD and LSUI datasets.
+The some examples are randomly selected on the UIEBD and LSUI datasets.
+Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from real scenes in Figure 6.
 Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
-4.3 Ablation Study
-To evaluate the impact of each strategy on the diffusion model, we conduct an ablation study.
+4.3. Ablation Study
+To evaluate the impact of each strategy on the diffusion model, we conduct a comprehensive ablation study.
 Table 2 shows the ablation results on the LSUI dataset.
-Model A means that we just use the normal transformer structure.
-Compared with model A, model B, C and D achieve better results, suggesting that the knowledge of physics can help diffusion model to better restore image.
-Model B outperforms C, suggesting that the prior information of the transmission map is more important than the background light.
-Model E achieves the best performance, proving that our all contribution is effective for UIE tasks.
-Model E underperforms model F.
-The possible reason for this is that there is still a gap between the generated physics information (the transmission map and the background light) and the ideal true value.
-Therefore, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
-Table 2: Ablation study on LSUI dataset.
-$T$ and $B$ refer to transmission map and background light prior information, respectively.
-$MS$ means multi-scale operation in our designed MS-FFN.
-$GAP$ denotes that we use global average pooling for the extraction of physical prior information from both the generated transmission map and background light.
-[…]
-A $\times$ $\times$ $\times$ $\times$ 28.35 0.8943
-B $\checkmark$ $\times$ $\times$ $\checkmark$ 24.71 0.9193
-C $\times$ $\checkmark$ $\times$ $\checkmark$ 25.84 0.9077
-D $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 23.06 0.9298
-E $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 25.44 0.9177
-F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 22.15 0.9354
-5 Conclusion
+We use a full-reference assessment metric (SSIM) and perceptual assessment metric (FID) to evaluate the ablation study, respectively.
+Model A means that we just employ the most basic diffusion transformer structure.
+Firstly, we investigate the influence of the PPG branch on the diffusion process.
+Compared with model A, model B, C and D achieve better results, suggesting that the prior knowledge of physics can help diffusion model to better restore image.
+Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T^{c}$ plays a highly effective role in guiding the diffusion process.
+Model D outperforms C, suggesting that GBB is also useful to exploit more helpful the background light information for the diffusion process.
+Subsequently, we explore the impact of the INR branch on the performance of the diffusion model.
+Model E outperforms D, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models.
+Model F outperforms E, implying that the periodic spatial encoding can enhance the representation ability of INR branch.
+Subsequently, for models G and H, we respectively delve into the effects of our designed CAM and PPU on the model performance.
+Both models G and H yield improved outcomes, proving the significant role that CAM and PPU play in extracting more valuable feature-level physical information.
+Finally, Ours achieves the best performance, proving that our all contribution is effective for UIE tasks.
+Moreover, we also examine the applicability of PPU in other baseline models.
+Table 3 showcases the adaptability of PPU across different UIE methods, demonstrating its generalizability potential to exploit the beneficial feature-level physics information.
+Limitations.
+As illustrated in Figure 7, there remains a discernible discrepancy between the reconstructed images from the PPG branch and the corresponding input images, indicating an existing gap between the physical information generated by our PPG branch and the ideal or ground-truth physical representation.
+Consequently, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
+Table 3. Ablation study with the PPU on UIEBD.
+Method Unet DW-water Ushape
+w/o PPU(PSNR/SSIM) 17.41/0.7769 21.88/0.8194 21.25/0.8453
+w PPU(PSNR/SSIM) 19.64/0.8157 21.23/0.8331 21.89/0.8570
+5. Conclusion
 In this paper, we develop a novel UIE framework, namely PA-Diff.
@@
 To the best of our knowledge, it is the first the diffusion model based on physical perception in image enhancement tasks.
-Our designed physics prior generation branch is a plug-and-play universal module to produce physics prior information, which can be integrated into any deep learning framework.
+Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into any deep learning network.
 PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.

### v3

@@
 Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement
-Abstract.
+Abstract
 Underwater visuals undergo various complex degradations, inevitably influencing the efficiency of underwater vision tasks.
-Recently, diffusion models were employed to underwater image enhancement (UIE) tasks, and gained SOTA performance.
+Recently, diffusion models were employed to underwater image enhancement (UIE) tasks, and gained the best performance.
 However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
-In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploiting the knowledge of physics to guide the diffusion process.
-PA-Diff consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
+In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploit physical knowledge to guide the diffusion process.
+PA-Diff consists of the Physics Prior Generation (PPG) branch, the Implicit Neural Reconstruction (INR) branch, and the Physics-aware Diffusion Transformer (PDT) branch.
 Our designed PPG branch aims to produce the prior knowledge of physics.
 With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
-INR Branch can learn robust feature representations from diverse underwater image via implicit neural representation, which reduces the difficulty of restoration for PDT branch.
-Extensive experiments prove that our method achieves best performance on UIE tasks.
+INR Branch can learn robust feature representations from diverse underwater images via implicit neural representation, which reduces the difficulty of restoration for PDT branch.
+Extensive experiments demonstrate that our method achieves the best performance on UIE tasks.
 The code is available at https://github.com/chenydong/PA-Diff.
-Underwater image enhancement, Physics model, Diffusion model, Transformer
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer vision
-1. Introduction
-The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics(McMahon and Plaku, 2021) and tracking underwater objects(de Langis and Sattar, 2020).
-Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness (Akkaynak et al., 2017).
-Consequently, the primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images.
-To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented (Peng and Cosman, 2017; Jr.
-et al., 2013; Peng et al., 2018).
+I Introduction
+The primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images.
+UIE stands as an indispensable component for tasks related to vision in the underwater domain.
+In recent years, UIE has been intensively studied, and many methods are proposed [29][40][64].
+Mathematically, most of these methods are characterized by the simplified atmospheric model:
+$I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$ (1)
+where $c\in\{R,G,B\}$ is the color channel, $I^{c}(x)$ is the underwater image, $J^{c}(x)$ is the scene direct reflected light at $x$, $t^{c}(x)$ is the medium transmission map, $B^{c}$ is the global background scattered light.
+The goal of the UIE is to recover the clear $J^{c}(x)$ from a given underwater image $I^{c}(x)$.
+Conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented [35, 18, 34].
 These techniques delve into the physical aspects contributing to degradation, such as color cast or scattering, seeking to rectify and enhance the underwater images.
-Nevertheless, these models grounded in physics exhibit restricted representational capacities, rendering them insufficient for comprehensively addressing the intricate physical and optical elements inherent in underwater scenes.
-Consequently, their efficacy diminishes, yielding suboptimal enhancement outcomes, particularly in the face of highly intricate and diverse underwater scenarios.
-In recent times, many learning-based methodologies (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fabbri et al., 2018) have been introduced to yield superior results.
+Nevertheless, these models grounded in physics exhibit restricted representational capacities, particularly in the face of highly intricate and diverse underwater scenarios.
+To address this problem, many learning-based methodologies [40, 33, 21, 10] have been introduced to yield superior results.
 Leveraging the potent feature representation and nonlinear mapping capabilities of neural networks, these methods excel.
-Deep learning-based methods can be roughly categorized into two classes: physics-free end-to-end methods (Jiang et al., 2022; Peng et al., 2023) and physics-aware methods (Li et al., 2021; Mu et al., 2023b; Fu et al., 2022a).
-Regarding the physics-free end-to-end methods, most of them usually use ground-truth images as labels to enforce L1/L2 distance-based consistency and also introduce various regularizations as additional constraints to deal with the ill-posed property.
-Although these methods achieve decent performance, they mainly consider the model in the raw image space, which does not fully explore the useful physics-aware feature information for UIE.
+Those methods can be roughly categorized into two classes: physics-free end-to-end methods [16, 33] and physics-aware methods [20, 29, 11].
+Regarding physics-free end-to-end methods, although they achieve decent performance, they mainly operate in the raw image space, which does not fully explore the useful physics-aware feature information for UIE.
+Moreover, the interpretability of physics-free models remains underexplored regarding the physical process of UIE.
 Regarding the physics-aware deep methods, most of them utilize the atmospheric scattering model.
-Existing UIE approaches (Li et al., 2021; Fu et al., 2022a) develop deep neural networks to estimate the transmission and the background scattered light, and then follow the traditional UIE methods to restore underwater images.
-However, these approaches do not correct the errors of the atmospheric light estimated by neural networks, which seriously affects the quality of the images (Ye et al., 2022).
+Existing UIE approaches [20, 11] develop deep neural networks to estimate the transmission and the background scattered light, and then follow the traditional UIE methods to restore underwater images.
+However, these approaches do not correct the errors of the atmospheric light estimated by neural networks, which seriously affects the quality of the images [55].
 Therefore, there is a crucial problem to adequately exploit the physical information and effectively incorporate them into a unified UIE deep framework.
-Recently, there has been a surge of interest in image synthesis (Rombach et al., 2022; Saharia et al., 2022b; Lu et al., 2024, 2023; Zhou et al., 2024) and restoration tasks (Choi et al., 2021; Wang et al., 2023; Yi et al., 2023; Zhou et al., 2023), with a notable focus on diffusion-based techniques like DDPM (Ho et al., 2020) and DDIM (Song et al., 2021).
-This heightened attention can be attributed to the remarkable generative capabilities of diffusion models.These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process (Ho et al., 2020), which poses a new perspective for underwater image enhancement task.
+Recently, there has been a surge of interest in image synthesis [24, 23, 71] and restoration tasks [6, 44, 56, 72], with a notable focus on diffusion-based techniques like DDPM [14] and DDIM [39].
+This heightened attention can be attributed to the remarkable generative capabilities of diffusion models.
 Tang et al.
-(Tang et al., 2023) present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
-UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023).
+[40] present an underwater image enhancement approach with diffusion model, WF-Diff [64] proposed a frequency diffusion model for UIE, and UIEDP proposed a UIE framework with Diffusion Prior[9].
 However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity for the diffusion model.
 Moreover, the diffusion models lack awareness regarding the difficulty of enhancing different image regions, a critical consideration for modeling the complex distribution in real-world underwater scenes.
+As shown in Figure 1, compared to previous diffusion-based method [40], our approach achieves superior generalization capability in real-world scenarios.
+Figure 1: Visualization comparison of 1080P images in real-world scenarios.
+Please zoom to view.
+Compared to DM-Water [40], both our full model and its version without diffusion technique exhibit stronger generalization capability for real-world underwater image enhancement, demonstrating the crucial role of physical priors in improving model generalization.
+Moreover, with the diffusion technique, our PA-Diff produces clearer results, further validating the effectiveness of the diffusion model.
 In this paper, we develop a novel physics-aware diffusion model, fully exploiting physical information to guide the diffusion process, called PA-Diff.
 It mainly consists of three branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
-The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model (Akkaynak and Treibitz, 2019a).
+The PPG branch aims to generate the transmission map and the global background light as physics prior information.
 The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability.
 The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch.
-Inspired by the latest works (Chen et al., 2024; Yang et al., 2023a) based on INR that can encode an image as a continuous function and learn a more robust degradation representation, we further incorporate the INR into our proposed framework to learn robust feature representations from diverse underwater image.
-The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information.
-Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA), physics perception unit (PPU) and gated multi-scale feed-forward network (GM-FFN) to exploit physics prior information and capture the long-range diffusion dependencies.
-Our PA-Diff achieves unprecedented perceptual performance in UIE task.
-Extensive experiments demonstrate that our developed PA-Diff performs the superiority against previous UIE approaches, and ablation study can demonstrate the effectiveness of all contributions.
+Inspired by the latest works [3, 52] based on INR that can encode an image as a continuous function and learn a more robust degradation representation, we further incorporate the INR into our proposed framework to learn robust feature representations from diverse underwater images.
+Finally, the PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information.
+Specifically, we design a physics-aware diffusion transformer block which contains a physics-aware self-attention (PA-SA), physics perception unit (PPU) and gated multi-scale feed-forward network (GM-FFN) to exploit physics prior information and capture the long-range diffusion dependencies.
+Our PA-Diff achieves unprecedented perceptual performance in UIE tasks.
+Extensive experiments demonstrate that our developed PA-Diff performs the superiority against previous UIE approaches, and comprehensive ablation studies validate the effectiveness of each proposed component.
 In summary, the main contributions of our PA-Diff are as follows:
   • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) branch, Implicit Neural Reconstruction (INR) branch, and Physics-aware Diffusion Transformer (PDT) branch.
-To the best of our knowledge, it is the first the Diffusion model with the guidance of physical knowledge in underwater image enhancement tasks.
-  • We incorporate Implicit Neural Reconstruction into our proposed PA-Diff to learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
+To the best of our knowledge, it is the first diffusion model with the guidance of physical knowledge in UIE tasks.
+  • We incorporate Implicit Neural Reconstruction (INR) into PA-Diff to learn robust feature representations from diverse underwater images, thereby reducing the difficulty of restoration for diffusion models.
   • We design a physics-aware diffusion transformer block, which not only enables PA-Diff with underwater-aware ability to guide the diffusion process, but also captures the long-range diffusion dependencies.
-Moreover, we also propose a physics perception unit (PPU), which can explore the beneficial feature-level physics information for underwater image enhancement.
-  • Extensive experiments compared with SOTAs considerably demonstrate that our PA-Diff performs the superiority against previous UIE approaches, and extensive ablation experiments can demonstrate the effectiveness of all contributions.
-Figure 1.
-Overall framework of our proposed PA-Diff.
-PA-Diff mainly consists of three cooperative branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
-Our designed PPG branch aims to produce the prior knowledge of physics.
-With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
-INR Branch can learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
-2. Related Works
-2.1. Underwater Image Enhancement
-Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches (Peng and Cosman, 2017; Peng et al., 2018; Tang et al., 2023; Peng et al., 2023; Li et al., 2020).
-Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors (Peng and Cosman, 2017), attenuation curve priors (Wang et al., 2018), fuzzy priors (Chiang and Chen, 2012).
-In addition, Akkaynak and Treibitz (Akkaynak and Treibitz, 2019b) proposed a method based on the revised physical imaging model.
-However, the depth map of the underwater scene is difficult to obtain.
-This leads to unstable performance of the this method, which usually suffers from severe color cast and artifacts.
-Therefore, the manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances.
-In recent years, underwater physical imaging models are gradually utilized in combination with data-driven methods (Li et al., 2021).
-Recently, deep learning-based methods (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fu et al., 2022b) achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance (Peng and Cosman, 2017; Jr.
-et al., 2013; Zhao et al., 2023).
-The previous methods neglect the underwater imaging mechanism and rely only on the representation ability of deep networks.
-Ucolor (Li et al., 2021) combined the underwater physical imaging model in the raw space and designed a medium transmission guided model.
-ATDCnet (Mu et al., 2023a) attempted to use transmission maps to guide deep neural networks.
+Moreover, we propose a Physics Perception Unit (PPU) to leverage beneficial feature-level physical priors for underwater image enhancement.
+  • Extensive experiments against best methods demonstrate that our PA-Diff consistently outperforms previous UIE approaches, and extensive ablation experiments can demonstrate the effectiveness of all contributions.
+The remainder of this paper is organized as follows.
+Section II presents the related work, followed by a detailed description of our PA-Diff in Section III.
+Experimental results are provided in Section IV, and conclusions are drawn in Section V.
+Figure 2: The overall framework of the proposed PA-Diff.
+A-Diff consists of three cooperative branches: the Physics Prior Generation (PPG) branch, the Implicit Neural Reconstruction (INR) branch, and the Physics-aware Diffusion Transformer (PDT) branch.
+The PPG branch aims to produce the prior knowledge of physics.
+With utilizing this prior to guide the diffusion process, the PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
+INR Branch can learn robust feature representations from diverse underwater images, which reduces the difficulty of restoration for the diffusion models.
+II Related Work
+II-A Underwater Image Enhancement
+Currently, existing UIE methods can be briefly categorized into the physical and deep model-based approaches [35, 34, 40, 33, 21].
+Most UIE methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors [35], attenuation curve priors [43], fuzzy priors [5].
+However, the manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances.
+In recent years, underwater physical imaging models are gradually utilized in combination with data-driven methods [20].
+Recently, deep learning-based methods [40, 33, 21, 66, 12] achieve acceptable performance, and some complex frameworks are proposed and achieve state-of-the-art (SOTA) performance [35, 18, 65].
+Wang et al.
+[42] introduced UIE-Net, incorporating a pixel-level perturbation technique to mitigate subtle texture variations, thereby enhancing both convergence efficiency and learning precision.
+Li et al.
+[21] developed Water-Net, utilizing algorithms like white balance, histogram equalization, and gamma correction to create diverse inputs for underwater imagery, refining them via a feature transformation module and merging them with confidence maps from a gated fusion framework to achieve the final output.
+Qi et al.
+[36] separated the underwater image improvement process into color adjustment and visibility enhancement stages, formulating CCMSR-Net to extract both local attributes and global structures.
+Peng et al.
+[33] presented a U-shaped Transformer combined with a multiscale feature fusion component and a spatially aware global feature unit, explicitly enhancing attention toward color channels and spatial zones undergoing severe degradation.
+WaterFormer [47] introduces a global–local Transformer to alleviate detail loss and incorporates a learnable Environment Adaptor to adaptively handle diverse underwater conditions.
+SSD-UIE [49] presents a semi-supervised domain-adaptive framework to mitigate both inter-domain and intra-domain discrepancies.
+The previous methods [33, 21, 18, 65] neglect the underwater imaging mechanism and rely only on the representation ability of deep networks.
+Ucolor [20] combined the underwater physical imaging model in the raw space and designed a medium transmission guided model.
+ATDCnet [28] attempted to use transmission maps to guide deep neural networks.
+SyreaNet [48] adopts a physically guided generative adversarial framework to address the generalization challenge in real underwater image enhancement.
+UIR-ES [74] proposes a physics-inspired unsupervised framework to improve data-efficient learning and enhance generalization in real-world scenarios.
 Therefore, how to fully exploit the knowledge of physics in deep neural networks is a very crucial problem.
-2.2. Diffusion Model
-Diffusion Probabilistic Models (DPMs) (Ho et al., 2020; Song et al., 2021) have been widely adopted for conditional image generation (Choi et al., 2021; Wang et al., 2023; Lu et al., 2023; Zhou et al., 2023, 2024).
-Saharia et al.
-(Saharia et al., 2022a) proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration.
-Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration (Özdenizci and Legenstein, 2023), document enhancement(Yang et al., 2023b), and low light image enhancement (Yang et al., 2023c; Zhou et al., 2023).
-Recently, Tang et al.
-(Tang et al., 2023) presented an image enhancement approach with diffusion model in underwater scenes.
-WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models.
-UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023).
+We aim to explore physics-aware diffusion framework, where physical priors are explicitly embedded into the diffusion process.
+II-B Diffusion Model
+The fusion of deep learning with nonequilibrium thermodynamics has driven diffusion model advancements.
+Through denoising, the network extracts key data features by recovering original information from noise.
+Thus, the diffusion model not only reconstructs data structure but also forms a trainable generation framework [63].
+Diffusion Probabilistic Models (DPMs) [14, 39] have been widely adopted for conditional image generation [44, 23, 72, 71, 68].
+Currently, diffusion models have been applied to a variety of low-level visual tasks such as image restoration [30], document enhancement[54], and low light image enhancement [53, 72].
+The diffusion model provides benefits for handling intricate noise conditions in underwater visuals, owing to its strong noise removal ability and clear interpretability [63].
+Tang et al.
+[40] present an underwater image enhancement approach with diffusion model and WF-Diff [64] proposed a frequency diffusion model for UIE.
+Lu et al.
+[25] later presented an approach for underwater image improvement utilizing denoising diffusion probabilistic models (UW-DDPM).
+They also developed a real-time UIE strategy founded on the denoising diffusion probabilistic framework: SU-DDPM [26].
+Presently, some scholars are exploring diffusion models to refine existing enhancement techniques.
+Du et al.
+[9] introduced UIE with diffusion prior (UIEDP), an innovative design that frames UIE as a posterior distribution sampling process, aiming to restore clear images from degraded underwater inputs.
+This method employs synthetic diffusion models to steer modern UIE algorithms, leading to enhanced results.
 However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
-2.3. Implicit Neural Representation
-Implicit Neural Representations (INRs) have emerged as an innovative technique garnering substantial attention for their ability to utilize coordinate-based Multi-Layer Perceptrons (MLPs) to model continuous-domain signals.
-Recently, some works have delved into the potential of INRs within the realm of two-dimensional image processing, including applications like image compression (Strümpler et al., 2022), and super-resolution(Chen et al., 2023).
-Chen et al.
-(Chen et al., 2021) first capitalized on INRs to achieve continuous image super-resolution.
-Subsequently, Lee et al.
-(Lee et al., 2022) proposed a dominant-frequency estimator that effectively predicts local texture information in natural images.
-Sun et al.
-(Sun et al., 2021) leveraged INRs to forecast continuous attributes based on the captured tomographic features.
-More recently, Quan et al.
-(Quan et al., 2023) introduced an INR-based inverse kernel prediction network to address the problem of single image defocus deblurring.
-Concurrently, Yang et al.
-(Yang et al., 2023a) exploited the controllable fitting capabilities of INRs to tackle the challenge of low-light image enhancement.
-Chen et al.
-(Chen et al., 2024) utilized INRs to learn common rain degradation representations from diverse rainy scenes.
-These efforts (Sun et al., 2021; Yang et al., 2023a; Chen et al., 2024) convincingly demonstrate the robust representation capabilities of INRs when applied to degraded images.
-Thus, inspired by the previous works, we explore the application of INRs to obtain robust feature representations for underwater image enhancement, aiming to reduce the difficulty of restoration for diffusion models.
-Figure 2. The detailed architecture of our designed (a) physics-aware self-attention (PA-SA) and (b) gated multi-scale feed-forward network (GM-FFN).
-3. Methodology
-3.1. Overall Framework
-The overall framework of PA-Diff is shown in Fig.
-1.
-PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch.
+III Methodology
+III-A Overall Framework
+The overall framework of PA-Diff is shown in Figure 2.
+PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which which mainly consists of the Physics Prior Generation (PPG) branch, the Implicit Neural Reconstruction (INR) branch, and the Physics-aware Diffusion Transformer (PDT) branch.
 The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model.
@@
 The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch.
-Inspired by (Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce implicit neural representations to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models.
+Inspired by [52], based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce INR to learn robust degradation representations from diverse underwater images, aiming to reduce the difficulty of restoration for diffusion models.
 The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information.
-We adopt the diffusion process proposed in DDPM (Ho et al., 2020) to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
+We adopt the diffusion process proposed in DDPM [14] to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
 Forward Diffusion Process.
@@
 Given a clean data $x_{0}$, then introduce Gaussian noise based on the time step, as follows:
-(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
+$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (2)
 where $\beta_{t}$ is a variable controlling the variance of the noise. Introducing $\alpha_{t}=1-\beta_{t}$, this process can be described as:
-(2) $x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).$
+$x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).$ (3)
 With Gaussian distributions are merged, we can obtain :
-(3) $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$
+$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$ (4)
 Reverse Diffusion Process. The reverse diffusion process aims to restore the clean data from the Gaussian noise. The reverse diffusion can be expressed as:
-(4) $p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$
-where $x_{c}$ refers to the conditional image ${I}$ (underwater input image).
-$\boldsymbol{\mu}_{\theta}(x_{t},x_{c},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively.
-We follow the setup of (Ho et al., 2020), they can be expressed as:
-(5) $\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$
-(6) $\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$
+$p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},x_{p},t),\sigma_{t}^{2}\mathcal{Z}),$ (5)
+where $x_{c}$ refers to the conditional image I{I}, and $x_{p}$ refers to the physics prior.
+$\boldsymbol{\mu}_{\theta}(x_{t},x_{c},x_{p},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively.
+We follow the setup of [14], they can be expressed as:
+$\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$ (6)
+$\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$ (7)
 where $\epsilon_{\theta}(x_{t},x_{c},t)$ is the estimated value with a Unet.
 We optimize an objective function for the noise estimated by the network and the noise $\epsilon$ actually added. Therefore, The diffusion loss is:
-(7) $L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$
-3.2. Physics Prior Generation
-Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model (Akkaynak and Treibitz, 2019a):
-(8) $I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$
-where $c\in\{R,G,B\}$ is the color channel, $x$ indicates the spatial location of each pixel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}$ is the medium transmission map, and $B^{c}$ is the global background scattered light.
-In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
-Network architecture of PPG branch is shown in Fig.
-1, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$.
-The two sub-networks consist of duplicated Dynamic Convolutions (DC) (Chen et al., 2020).
+$L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},x_{p},t)\|.$ (8)
+III-B Physics Prior Generation
+Mathematically, the underwater images can be expressed by the simplified Koschmieder light scanning model [1]:
+$I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$ (9)
+where $c\in\{R,G,B\}$ is the color channel, $I^{c}(x)$ is the observed underwater image intensity at pixel $x$ in channel $c$, consisting of both attenuated scene radiance and veiling-light backscatter; $J^{c}(x)$ is the latent clean radiance (clear-water equivalent), representing the true scene color if the medium were air instead of water; $T^{c}(x)$ is the transmission map, indicating the fraction of light that reaches the camera without being absorbed or scattered, depending on the scene depth and the wavelength-dependent attenuation coefficient; $B^{c}$ is the background light, modeling ambient light scattered into the camera from distant water volumes where object visibility diminishes.
+In the PPG branch, our goal is to produce $T^{c}(x)$ and $B^{c}$ from the $I^{c}(x)$.
+Network architecture of PPG branch is shown in Figure 2, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$.
+The two sub-networks mainly consist of duplicated Dynamic Convolutions (DC) [4].
 DC dynamically combines numerous parallel convolutional kernels according to their attention, thereby enhancing the model’s complexity and performance without augmenting the network’s depth or breadth.
-Given that the background light differs fundamentally from the transmission map, being independent of image content and instead encapsulating the global properties (Fu et al., 2022a) , for the purpose of generating the background light component, our background light generation sub-network initially employs a Gaussian blur block (GBB) to eliminate fine-grained content details and thereby extract the inherent global attributes.
-Given a underwater image as input $I$ and its corresponding ground truth (GT), we can obtain via Eq. 8:
-(9) $\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(GBB(I)),$
+Moreover, we introduce channel and spatial attention block to enhance the ability of feature learning.
+Given that the background light differs from the transmission map and is independent of image content, it instead encapsulates global properties [11].
+For the purpose of generating the background light component, our background light generation sub-network initially employs a Gaussian blur block (GBB) to eliminate fine-grained content details and thereby extract the inherent global attributes.
+Note that, for the background light generation sub-network, we first perform gaussian blur module to remove content details and obtain the global attribute.
+Different from the transmission map, the global background light is independent of the image content and reflects the global property.
+Given an underwater image as input $I$ and its corresponding ground truth (GT), we can obtain via Eq. 7:
+$\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(GBB(I)),$ (10)
 We constrain the PPG branch by supervised manner. The physics reconstruction loss of PPG branch can be expressed as:
-(10) $\mathcal{L}_{p}=||\hat{I}-I||_{1}.$
+[…]
 To enhance the perceptual similarity of the reconstructed images, we introduce the perceptual Loss $\mathcal{L}_{per}$ :
-(11) $\mathcal{L}_{per}=||\phi(\hat{I})-\phi(I)||_{2},$
+[…]
 where $\phi$ is a pre-trained VGG network.
-Finally, the overall loss $\mathcal{L}_{ppg}$ of PPG branch can be expressed as:
-(12) $\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$
-Figure 3. The detailed structure of the proposed physics perception unit (PPU).
-3.3. Implicit Neural Reconstruction
-Inspired by (Quan et al., 2023; Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models.
-INRs trains a multi-layer perceptron (MLP) by learning the following mapping :
-(13) $F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$
-where the dimensions of input refer to the (x, y) spatial coordinates, while the dimensions of output refer to the (R, G, B) color channels.
-Specifically, following the setting of (Chen et al., 2024; Yang et al., 2023a), underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with a image resolution of H × W.
+Finally, the overall loss $\mathcal{L}_{ppg}$ of the PPG branch can be expressed as:
+$\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$ (13)
+Figure 3: The detailed architecture of our designed (a) physics-aware self-attention (PA-SA) and (b) gated multi-scale feed-forward network (GM-FFN).
+Figure 4: The detailed structure of the proposed physics perception unit (PPU).
+To the best of our knowledge, it is the first method which is based on the interpretable physics model in the feature space for challenging underwater image enhancement tasks.
+III-C Implicit Neural Reconstruction
+Inspired by [37, 3, 52] based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater images, aiming to reduce the difficulty of restoration for diffusion models.
+INRs train a multi-layer perceptron (MLP) by learning the following mapping :
+$F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$ (14)
+$\mathbb{R}^{2}$ denotes the continuous 2D spatial coordinate domain $(x,y)$ and $\mathbb{R}^{3}$ denotes the corresponding RGB color space, i.e., each spatial location is mapped to its RGB color value.
+Specifically, following the setting of [3, 52], an underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with an image resolution of H × W.
 Meanwhile, the position of each pixel is recorded in a relative coordinate set $\mathbf{X}\in\mathbb{R}^{H\times W\times 2}$, where the value ’2’ means horizontal and vertical coordinates.
 As shown in Figure 2, we fuse $X$ and $E$, and use $F_{\theta}$ to output image $\boldsymbol{I}_{{\mathrm{inr}}}$, which is expressed as:
-(14) $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$
+$\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$ (15)
 where $[x,y]$ is the location of a pixel and $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]$ is the generated RGB value.
-Due to the elimination of high-frequency components during rendering (Yang et al., 2023a), we adopt periodic spatial encoding (Lee and Jin, 2022) to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details.
+Due to the elimination of high-frequency components during rendering [52], we adopt periodic spatial encoding [19] to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details.
 The encoding procedure is formulated as:
-(15) $\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$
-(16) $\gamma(c)=(\cdot\cdot\cdot,\sin(2^{i}c),\cos(2^{i}c),\cdot\cdot\cdot),$
+$\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$ (16)
+$\gamma(x)=(\cdot\cdot\cdot,\sin(2^{i}x),\cos(2^{i}x),\cdot\cdot\cdot),$ (17)
 where $\gamma()$ represents a spatial encoding function.
-$i$ values from 00 to $L-1$, $c\in\mathbb{R}^{2L}$ is the coordinate value of $X$, and $L$ is a hyperparameter that determines the dimension value.
+$i$ values from 0 to L1L-1, and $L$ is a hyperparameter that determines the dimension value.
 Given a GT of underwater image $GT$, The reconstruction loss of INR branch can be expressed as:
-(17) $\mathcal{L}_{INR}=||\boldsymbol{I}_{\mathrm{inr}}-GT||_{1}.$
+[…]
 Finally, we fuse $I$ and $\boldsymbol{I}_{\mathrm{inr}}$ to obtain the condition $x_{c}$ of diffusion model, which can be expressed as:
-(18) $x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$
-3.4. Physics-aware Diffusion Transformer
+$x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$ (19)
+III-D Physics-aware Diffusion Transformer
 Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration.
-Fig.
-1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention module (CAM).
-Given the output of the INR branch as condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch.
+Figure 5 showcases the prior results provided by our PPG branch and INR branch.
+Figure 2 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention modules (CAM).
+Given the output of the INR branch as the condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch.
 Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel.
@@
 We denote $T^{c}$ and $\mathbf{F}$ as the input of CAM.
-We use different linear projections to construct $Q$ and $K$ in CAM:
-(19) $Q=Conv_{1\times 1}(T^{c}),$
-(20) $K=Conv_{1\times 1}(\mathbf{F}).$
-Similarly, $V$ of the transmission map $T^{c}$ can be obtained:
-(21) $V=Conv_{1\times 1}(T^{c}).$
-The output feature map of CAM $T_{out}$ can then be obtained from the formula:
-(22) $T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$
-Physics-aware Self-attention. As is shown in Figure 2 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA:
-(23) $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t},$
-where $Norm$ means layer normalization.
-We first obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$.
+We use different linear projections to construct $Q$ and $K$ in CAM: $Q=Conv_{1\times 1}(T^{c}),K=Conv_{1\times 1}(\mathbf{F}).$ Similarly, $V$ of the transmission map $T^{c}$ can be obtained: $V=Conv_{1\times 1}(T^{c}).$ The output feature map of CAM $T_{out}$ can then be obtained from the formula:
+$T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$ (20)
+Physics-aware Self-attention.
+As is shown in Figure 3 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA: $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t}$.
+We obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$.
 Meanwhile, we generate query ($Q_{t}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{T_{out}}$) from the output feature map of CAM, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution.
 We employ physics-aware query-driven attention strategy to obtain the final query $Q$ by fusing $Q_{t}$ and $Q_{f}$, which can be expressed as:
-(24) $Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$
+$Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$ (21)
 We employ self-attention and get the output of PA-SA $\mathbf{\hat{F}}$:
-(25) $\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$
+$\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$ (22)
 where $\alpha$ is a learnable parameter.
@@
 Physics Perception Unit.
-Exploring feature-level physics information (Zheng et al., 2023; Dong and Pan, 2020) achieved success in image dehazing tasks.
+Exploring feature-level physics information [70, 8] achieved success in image dehazing tasks.
 Our proposed physics perception unit (PPU) aims to further exploit the beneficial feature-level physics information for underwater image enhancement, and learn richer physical features.
 Our proposed PPU is based on the simplified atmospheric model Eq.
-(8), as shown in Fig.
-3.
+(5), as shown in Figure 4.
 We can reformulate the physics model to represent the clear image J, which can be obtained by:
-(26) $J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$
+$J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$ (23)
 Let $k$ denote a feature extractor(the filter kernel in a deep network). Then extracting features via the feature extractor k, Eq. (17) can be reformulated as follows:
-(27) $k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$
+$k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$ (24)
 where $\otimes$ denotes the convolution operator and $\odot$ denotes the element-wise product operation. Using the matrix-vector form for algebraic operations, Eq. (26) can be express as:
-(28) $\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$
+$\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$ (25)
 where $\boldsymbol{K},\boldsymbol{J},\boldsymbol{I},\boldsymbol{B},\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ respectively denote the matrix-vector forms of $k,J,I,B,\frac{1}{t}$ and $(1-\frac{1}{t})$.
 Note that the diagonal vector of the diagonal matrix $\boldsymbol{T}_{1}$ corresponds to the vectorized form of $\frac{1}{t}$.
-Assuming that $\boldsymbol{K}$ is an invertible matrix-vector, we can obtain
-(29) $\boldsymbol{F}_{1}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1},$
-(30) $\boldsymbol{F}_{2}=\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1}.$
-And then, Eq. (27) can be express as:
-(31) $\boldsymbol{KJ}$ $=(\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{I}+(\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{B}$
-$=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$
+Then, we can decompose the matrix $\boldsymbol{K}\boldsymbol{T}_{1}$ and $\boldsymbol{K}\boldsymbol{T}_{2}$ into product of two matrices $\boldsymbol{F}_{1}\boldsymbol{K}$ and $\boldsymbol{F}_{2}\boldsymbol{K}$.
+Eq.
+(19) can be express as:
+$\boldsymbol{KJ}$ $=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$ (26)
 The Eq.
@@
 (20), the $\tilde{\boldsymbol{J}}$ can be express as:
-(32) $\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}.$
+$\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}$. (27)
 Since the background light is usually assumed to be homogeneous, we can use global average pooling (GAP) to obtain the main information for $\tilde{\boldsymbol{B}}$ in the feature space.
 $\tilde{\boldsymbol{B}}$ can be express:
-(33) $\tilde{\boldsymbol{B}}=H(\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{GAP}(\tilde{\boldsymbol{I}}))))))))),$
+[…]
 where $\sigma$ is the Sigmoid function, H denotes a replication operation.
 As the transmission map is non-homogeneous, we cannot apply GAP for the approximations of $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$.
-We use different projection networks to obtain $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ :
-(34) $\tilde{\boldsymbol{t}_{1}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))),$
-(35) $\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))).$
-where s represents the shared Conv layer. Given $\hat{\mathbf{F}}$ as the input of PPU, we use residual connection to obtain the final output of PPU, which can be expressed as :
-(36) $\hat{\mathbf{F_{ppu}}}=PPU(\hat{\mathbf{F}})+\hat{\mathbf{F}}.$
+$\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ can be obtained by
+$))),\tilde{\boldsymbol{t}_{1}},\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv}(\tilde{\boldsymbol{I}}$)))))), (29)
 Gated Multi-scale Feed-forward Network.
-Finaly, we design a GM-FFN for local feature aggregation.
-As is shown in Figure 2 (b), we employ GD-FFN-like structure (Zamir et al., 2022) to construct our feed-forward network.
+Finally, we design a GM-FFN for local feature aggregation.
+As is shown in Figure 3 (b), we employ GD-FFN-like structure [57] to construct our feed-forward network.
 In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions.
 GM-FFN adopts GELU to ensure the flexibility of feature aggregation.
-Figure 4. The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
-Table 1. Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
-Methods UIEWD UWCNN UIECˆ2-Net Water-Net U-color U-shape DM-water Ours
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 5. Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
-Figure 6. More visual results from real-world scenarios.
-Figure 7. Limitations. The output images refer to the reconstruction images of PPG branch.
-4. Experiments
-4.1. Experimental Settings
+Discussion on Prior Multi-scale Works.
+MB-TaylorFormerv2 [17] introduces multi-scale patch embedding in transformer for image restoration, incorporating varying receptive field sizes, flexible receptive field shapes, and multi-level semantic information.
+ESTINet [58] proposes a feature-level multi-scale fusion module between the encoder and decoder for video deraining.
+DDMSNet [60] presents a multi-stage coarse-to-fine multi-scale CNN framework for snow removal, achieving progressive restoration.
+GridFormer [41] adopts a pixel-level multi-scale transformer architecture for image restoration, effectively leveraging complementary multi-scale information in the pixel space.
+Other works [61, 59, 67, 38, 69] also explore multi-scale information fusion and utilization at both the pixel and feature levels.
+In contrast, our work investigates multi-scale information utilization within the feed-forward network (FFN) of transformers, enhancing the FFN’s ability to capture multi-scale features.
+Figure 5: The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
+Figure 6: Qualitative comparison with other SOTA methods on the LSUI dataset.
+(a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
+Figure 7: Qualitative comparison with other SOTA methods on the UIEB dataset.
+(a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
+TABLE I: Implementation details of our PA-Diff.
+Hyperparameter Value
+Optimizer Adam
+Learning Rate $1\times 10^{-4}$ (fixed)
+Batch Size 6
+Total Iterations 1,050,000
+Input Resolution $256\times 256$
+Diffusion Timesteps 2,000
+Beta Schedule Linear ($1\times 10^{-6}$ to $1\times 10^{-2}$)
+UNet Inner Channel 48
+Channel Multiplier $[1,2,4,8,8]$
+Attention Resolution 16
+Dropout 0.2
+Norm Groups 24
+GPU Hardware NVIDIA RTX 3090 (24GB)
+Training Time per Epoch $\sim$8.6 minutes
+Training Time per Iteration $\sim$0.8 seconds
+TABLE II: Quantitative comparison on the UIEB and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
+Method UIEB LSUI
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
+UIEWD [27] 13.68 0.696 0.437 174.129 0.607 4.132 14.45 0.658 0.429 101.326 0.547 4.369
+UWCNN [2] 13.94 0.716 0.335 91.048 0.551 4.082 16.59 0.706 0.361 61.202 0.551 4.103
+WaterNet [33] 22.76 0.906 0.129 42.432 0.615 4.341 20.78 0.804 0.234 37.900 0.569 4.349
+UIEC2Net [45] 20.04 0.818 0.178 49.071 0.606 4.369 24.25 0.854 0.142 37.425 0.591 4.307
+Ushape [21] 18.15 0.833 0.210 68.359 0.555 4.235 23.66 0.845 0.158 29.051 0.573 4.351
+Ucolor [20] 20.08 0.830 0.184 73.245 0.587 4.236 23.98 0.864 0.166 35.279 0.566 4.235
+GUPDM [29] 19.90 0.877 0.157 41.825 0.573 4.226 24.48 0.869 0.119 29.360 0.578 4.380
+CECF [7] 22.20 0.841 0.135 32.370 0.619 4.374 26.58 0.886 0.115 19.636 0.590 4.355
+DM-water [40] 22.78 0.894 0.113 29.249 0.598 4.217 26.95 0.881 0.143 21.433 0.589 4.257
+WF-Diff [64] 22.84 0.898 0.114 28.482 0.609 4.293 26.07 0.884 0.121 20.846 0.587 4.369
+HCLR [73] 22.87 0.919 0.112 27.864 0.607 4.292 26.33 0.880 0.108 27.211 0.593 4.408
+SS-UIE [32] 21.98 0.907 0.121 36.479 0.589 4.263 24.41 0.840 0.164 26.196 0.575 4.318
+PA-Diff 23.47 0.897 0.106 23.631 0.621 4.385 27.25 0.889 0.103 20.463 0.594 4.393
+TABLE III: Quantitative comparison on EUVP, U45, and Challenge subsets. The best results are highlighted in bold and the second best results are underlined.
+Method EUVP U45 Challenge
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$ UCIQE $\uparrow$ UIQM $\uparrow$ UCIQE $\uparrow$ UIQM $\uparrow$
+UIEWD [27] 13.93 0.623 0.452 91.712 0.504 4.063 0.510 4.057 0.550 3.641
+UWCNN [2] 17.50 0.692 0.341 55.767 0.570 4.157 0.534 4.075 0.535 3.671
+WaterNet [33] 18.43 0.770 0.308 39.888 0.585 4.182 0.552 4.055 0.579 3.839
+UIEC2Net [45] 24.04 0.818 0.180 29.755 0.594 4.248 0.575 4.105 0.581 3.935
+Ushape [21] 25.53 0.818 0.139 24.914 0.587 4.238 0.567 4.139 0.530 3.820
+Ucolor [20] 25.68 0.871 0.159 26.114 0.586 4.240 0.560 4.014 0.542 3.683
+GUPDM [29] 25.18 0.848 0.114 20.744 0.590 4.399 0.566 4.144 0.543 3.827
+CECF [7] 28.02 0.861 0.126 22.630 0.596 4.252 0.585 4.128 0.575 3.899
+DM-water [40] 26.45 0.852 0.183 24.484 0.598 4.266 0.592 4.139 0.586 3.919
+WF-Diff [40] 27.24 0.851 0.128 21.517 0.592 4.269 0.584 4.137 0.584 3.926
+HCLR [73] 27.08 0.850 0.117 25.114 0.591 4.288 0.584 4.117 0.562 3.792
+SS-UIE [32] 26.00 0.824 0.160 24.829 0.587 4.353 0.584 4.105 0.557 3.970
+PA-Diff 29.12 0.880 0.101 18.769 0.599 4.273 0.593 4.147 0.586 3.954
+Figure 8: Qualitative comparison with other SOTA methods on the Challenge dataset.
+(a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
+Figure 9: Visual analysis with ablation study on the LSUI datasets.
+Figure 10: Comparative analysis of diffusion-based underwater image restoration with and without physical priors.
+Figure 11: Multi-view visual comparison on UwMVS real-world test set [50].
+Figure 12: Limitations. The output images refer to the reconstruction images of PPG branch.
+Figure 13: Visual results under extreme low-light conditions.
+IV Experiments
+IV-A Experimental Settings
 Implementation Details.
@@
 The learning rate was established at 0.0001.
-During the training stage, to balance the batch size and image size, the batch and image size are set to 2 and 256 × 256, respectively.
+During the training stage, to balance the batch size and image size, the batch size and image resolution are set to 6 and 256 × 256, respectively.
 The number of training iterations reached one million.
@@
 The time step of the diffusion model is set to 2000.
-In the testing stage, following the setting in (Tang et al., 2023), the input size of the image is 256 × 256.
-By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
+A skip sampling strategy is employed with 10 sampling steps to balance enhancement performance and inference runtime.
+Table I shows the key details regarding the implementation of our PA-Diff.
 Datasets.
-We utilize the real-world UIEBD dataset (Peng et al., 2023) and the LSUI dataset (Li et al., 2020) for training and evaluating our model.
-The UIEBD dataset comprises 890 underwater images with corresponding labels.
-Out of these, 700 images are allocated for training, and the remaining 190 are designated for testing.
-LSUI dataset contains 5004 underwater images and their corresponding high-quality images.
-Compared with UIEBD, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images.
-In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing.
-In addition, to verify the generalization of PA-Diff, we use non-reference benchmarks U45 (Li et al., 2019), which contains 45 underwater images for testing.
+We utilize the UIEB dataset [21] and the LSUI dataset [33] for training and evaluating our model.
+The UIEB dataset comprises 890 underwater images with corresponding labels.
+Out of these, 800 images are allocated for training, and the remaining 90 are designated for testing.
+LSUI dataset contains 4279 underwater images and their corresponding high-quality images.
+Compared with UIEB, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images.
+In the paper, the LSUI dataset is randomly partitioned into 3879 images for training and 400 images for testing.
+In addition, to verify the generalization, we use the reference based benchmark EUVP [15], which contains 515 paired test samples, along with the non-reference based benchmarks U45 [22] and Challenge [21], which contain 45 and 60 underwater images, respectively.
 Comparison Methods.
-We conduct a comparative analysis between PA-Diff and seven state-of-the-art (SOTA) UIE methods, namely UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UWCNN (Anwar et al., 2018), UIEWD (Ma and Oh, 2022), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023).
+We conduct a comparative analysis with seven SOTA UIE methods, namely UIECˆ2-Net [45], Water-Net [33], UWCNN [2], UIEWD [27], U-color [20], U-shape [21], GUPDM [29], CECF [7], DM-water [40], WF-Diff [64], HCLR [73], and SS-UIE [32].
 To ensure a fair and rigorous comparison, we utilize the provided source codes from the respective authors and adhere strictly to the identical experimental settings across all evaluations.
 Evaluation Metrics.
-We employ standard full-reference image quality assessment metrics, including PSNR and SSIM (Wang et al., 2004), for quantitative comparisons at both pixel and structural levels.
-Higher PSNR and SSIM values indicate superior image quality.
-Additionally, LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) are utilized to evaluate perceptual performance.
-Lower LPIPS and FID scores signify a more effective UIE approach.
-For non-reference benchmarks U45, we introduce UIQM (Panetta et al., 2015) and UCIQE (Yang and Sowmya, 2015) to evaluate our method.
-Table 2.
-Ablation study on LSUI dataset.
-$T^{c}$ and $B^{c}$ refer to transmission map and background light prior information in PPG branch, respectively.
+We employ PSNR and SSIM [46] for quantitative comparisons at both pixel and structural levels.
+Additionally, LPIPS [62] and FID [13] are utilized to evaluate perceptual performance.
+For non-reference benchmarks U45 and Challenge, we introduce UIQM [31] and UCIQE [51] to evaluate our method.
+IV-B Results and Comparisons
+Table II shows the quantitative results compared with different SOTA methods on the UIEB and LSUI datasets.
+We mainly use PSNR, SSIM, LPIPS, FID, UCIQE and UIQM as our quantitative indices for UIEBD and LSUI datasets.
+The results show that our algorithm outperforms state-of-the art methods obviously, especially in terms of all metrics.
+Figure 6 and 7 present visual comparison results with other methods on the LSUI and UIEB datasets, with selected examples randomly chosen from each dataset.
+Furthermore, Table III presents a comparison using pre-trained models on LSUI across multiple datasets, demonstrating the superior generalization capability of our method.
+Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from the Challenge dataset in Figure 8.
+Our method consistently generates more natural and visually pleasing results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
+TABLE IV: Ablation with application of diffusion technology.
+Methods LSUI EUVP
+Name PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+w/o Diffusion 27.13 0.883 0.127 27.75 0.868 0.153
+PA-Diff (Ours) 27.25 0.889 0.103 29.12 0.880 0.101
+TABLE V: Ablation study on the LSUI dataset.
+$T_{c}$ and $B_{c}$ refer to transmission map and background light prior information in the PPG branch, respectively.
 GBB means Gaussian blur module.
 INR represents our INR branch.
-PSE refers to the periodic spatial encoding (Lee and Jin, 2022), aiming to better recover high-frequency details.
-CAM and PPU means our designed cross attention module and physics perception unit , respectively.
-$GM-FFN$ means multi-scale operation in the feed-forward network.
+PSE refers to the periodic spatial encoding [19], aiming to better recover high-frequency details.
+CAM and PPU denote the cross attention module and physics perception unit, respectively.
+GMFFNGM-FFN means multi-scale operation in the feed-forward network.
 […]
-A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 28.35 0.8893
-B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 27.71 0.8974
-C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.92 0.9071
-D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.24 0.9112
-E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 24.50 0.9186
-F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 24.92 0.9176
-G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 23.66 0.9241
-H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 22.85 0.9374
-Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 22.13 0.9378
-4.2. Results and Comparisons
-Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD, LSUI and U45 datasets, including with UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UIEWD (Ma and Oh, 2022), UWCNN (Anwar et al., 2018), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023).
-We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets, and UIQM and UCIQE for non-reference dataset U45.
-The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics.
-To better validate the superiority of our methods, Fig.
-5 shows the visual comparison results with other methods on the UIEBD and LSUI datasets.
-The some examples are randomly selected on the UIEBD and LSUI datasets.
-Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from real scenes in Figure 6.
-Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
-4.3. Ablation Study
+A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.14 0.852
+B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.67 0.857
+C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.73 0.865
+D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.94 0.868
+E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 26.47 0.873
+F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 26.51 0.872
+G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 26.79 0.880
+H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 27.04 0.883
+Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 27.25 0.889
+TABLE VI: Comparison results of inference speed and Params.
+Method WaterNet U-color Ushape DM-water PA-Diff
+Inference time(s) 0.61 2.75 0.03 0.14 0.23
+Params(M) 24.81 157.4 65.60 10.69 41.76
+TABLE VII: Params analysis of each branch in PA-Diff.
+Method PPG branch INR branch PDT branch PA-Diff (all)
+Params(M) 0.0302 0.1487 41.31 41.76
+IV-C Ablation Study
+Effects of Diffusion Technology.
+Firstly, to validate the value of the diffusion technique within our framework, Table IV presents the ablation study on the application of diffusion technology.
+The proposed model with diffusion technique achieves better results, demonstrating the significant performance improvements it brings to the model.
+Analysis of Model Complexity.
+Table VI shows a comparison of inference speeds and model parameters with the previous methods, indicating that our method demonstrates competitive model parameters.
+Table VII presents the network parameters for each branch, clearly indicating that the PPG and INR branches, as auxiliary networks for prior generation, have a minimal number of parameters and exhibit low complexity.
+However, we employ a transformer-based denoising network in the PDT branch, resulting in relatively large parameters.
+Therefore, our future endeavors will concentrate on investigating lighter-weight denoising networks.
+Impact of Different Framework.
 To evaluate the impact of each strategy on the diffusion model, we conduct a comprehensive ablation study.
-Table 2 shows the ablation results on the LSUI dataset.
-We use a full-reference assessment metric (SSIM) and perceptual assessment metric (FID) to evaluate the ablation study, respectively.
+Table V shows the ablation results on the LSUI dataset, and Figure 5 presents the visual results of the prior knowledge.
 Model A means that we just employ the most basic diffusion transformer structure.
 Firstly, we investigate the influence of the PPG branch on the diffusion process.
-Compared with model A, model B, C and D achieve better results, suggesting that the prior knowledge of physics can help diffusion model to better restore image.
-Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T^{c}$ plays a highly effective role in guiding the diffusion process.
-Model D outperforms C, suggesting that GBB is also useful to exploit more helpful the background light information for the diffusion process.
+Compared with model A, model B, C and D achieve better results, suggesting that physical prior knowledge helps the diffusion model better restore underwater images.
+Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T_{c}$ plays a highly effective role in guiding the diffusion process.
 Subsequently, we explore the impact of the INR branch on the performance of the diffusion model.
@@
 Model F outperforms E, implying that the periodic spatial encoding can enhance the representation ability of INR branch.
-Subsequently, for models G and H, we respectively delve into the effects of our designed CAM and PPU on the model performance.
-Both models G and H yield improved outcomes, proving the significant role that CAM and PPU play in extracting more valuable feature-level physical information.
-Finally, Ours achieves the best performance, proving that our all contribution is effective for UIE tasks.
-Moreover, we also examine the applicability of PPU in other baseline models.
-Table 3 showcases the adaptability of PPU across different UIE methods, demonstrating its generalizability potential to exploit the beneficial feature-level physics information.
+Finally, Ours achieves the best performance, proving that all our proposed components are effective for UIE tasks.
+Analysis and Discussion.
+We conduct an exhaustive analysis of the individual contributions of each core component to the holistic performance by visually presenting the results of ablation studies, as illustrated in Figure 9.
+Here, Model b represents the employment of the fundamental diffusion transformer architecture alone, serving as our baseline.
+Compared to the full model, the results of Model b show that the visual effects obtained by using only the basic diffusion model can be unrealistic.
+Models c and d are designed to investigate the influence of incorporating prior physical knowledge—embodied by the transmission map $T_{c}$ and background light information $B_{c}$, respectively—into the full model framework.
+Compared to model d, the results for model c are worse, mainly in terms of poorer recovery in some areas and less good colour recovery, suggesting that the transmission map $T_{c}$ we introduced plays a very important role.
+Model e, on the other hand, focuses on elucidating the impact of integrating an Implicit Neural Representation (INR) branch.
+Compared to the full model, the results of Model e show that the visual effects obtained by using only the basic diffusion model can be bad, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models.
+Lastly, Model F is pivotal in examining the role of the PPU module within the full model context.
+The visual evidence gathered unmistakably highlights the crucial role each segment of our proposed model plays in achieving optimal performance.
+Impact of Physical Priors.
+As illustrated in Figure 10, we show a comparative analysis of diffusion-based underwater image restoration with and without physical priors.
+The results demonstrate that, in the absence of physical priors, the diffusion model exhibits noticeable deviations in both global color correction and local detail reconstruction, particularly under severely degraded underwater conditions.
+In contrast, incorporating physical priors provides a robust guidance signal throughout the diffusion process, effectively alleviating error accumulation during reverse denoising.
+This design enables the diffusion model to better align with the underlying degradation mechanism, thereby improving its generalization capability and restoration robustness.
+The Plug-and-play Capability of PPU.
+Furthermore, we explore the application of the PPU in other models.
+Table VIII showcases the adaptability of PPU across different methods, demonstrating its universal potential across diverse methods to exploit the beneficial feature-level physics information.
+Impact of Different PPU Structure.
+We conduct an in-depth analysis of the key design choices in the proposed PPU.
+Specifically, we examine two critical physical parameters, $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$, which are closely related to the transmission map.
+As shown in Table IX, removing either parameter leads to a significant performance degradation, validating the effectiveness of the current PPU design.
+Furthermore, we include comparisons with alternative feature-level physics modeling modules, such as the PDU used in C2PNet [70].
+The results demonstrate that our PPU consistently achieves superior performance, indicating that the proposed PPU is more suitable for underwater image enhancement.
+Choice of Attention in the PPG Branch.
+We perform ablation studies on the specific attention mechanisms adopted in the PPG branch, as reported in Table X.
+Variants that remove either channel attention or spatial attention are evaluated.
+The results show a consistent, albeit slight, performance drop when either component is removed, supporting the effectiveness of our attention design.
+A Lightweight Variant of Our PA-Diff.
+we design a lightweight variant of our model by reducing the channel width and the number of modules in the transformer-based PDT.
+As reported in Table XI, the proposed lightweight PA-Diff (Mini) achieves performance comparable to the full model across all evaluation metrics, while using only about one quarter of the parameters of the full model.
+Moreover, when compared with DM-Water under a similar parameter budget, our lightweight model consistently delivers superior results, further validating the effectiveness and scalability of the proposed framework.
+Extention on Downstream Tasks.
+As shown in Figure 11, we provide a multi-view visual comparison using the UwMVS real-world test set [50], which further validate the effectiveness of our proposed method in improving reconstruction quality.
+TABLE VIII: Ablation study with the PPU on UIEBD.
+Method Unet DW-water Ushape
+w/o PPU(PSNR/SSIM) 17.41/0.7769 22.78/0.894 18.15/0.833
+w PPU(PSNR/SSIM) 19.64/0.8157 22.93/0.897 20.54/0.848
+TABLE IX: Ablation with the choice of PPU structure on the LSUI dataset.
+LSUI
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
+w/o $\tilde{\boldsymbol{t}_{1}}$ 26.31 0.880 0.130 24.61 0.590 4.306
+w/o $\tilde{\boldsymbol{t}_{2}}$ 25.88 0.876 0.135 33.96 0.585 4.232
+PDU [70] 26.94 0.878 0.126 22.76 0.593 4.318
+PA-Diff 27.25 0.889 0.103 20.46 0.594 4.393
+TABLE X: Ablation with the attention mechanisms of the PPG Branch on the LSUI dataset. CA and SA denote channel attention and spatial attention, respectively.
+LSUI
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
+w/o CA 27.11 0.885 0.110 22.14 0.592 4.372
+w/o SA 27.08 0.879 0.104 21.83 0.589 4.381
+PA-Diff 27.25 0.889 0.103 20.46 0.594 4.393
+TABLE XI: Comparison with lightweight variants on the LSUI dataset.
+LSUI
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ UIQM $\uparrow$ Params(M) $\downarrow$
+Ushape 23.66 0.845 0.158 4.351 65.60
+DM-water 26.95 0.881 0.143 4.257 10.69
+PA-Diff (Mini) 27.18 0.889 0.123 4.378 11.58
+PA-Diff (Ours) 27.25 0.889 0.103 4.393 41.76
 Limitations.
 As illustrated in Figure 12, there remains a discernible discrepancy between the reconstructed images from the PPG branch and the corresponding input images, indicating an existing gap between the physical information generated by our PPG branch and the ideal or ground-truth physical representation.
-Consequently, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
-Table 3. Ablation study with the PPU on UIEBD.
-Method Unet DW-water Ushape
-w/o PPU(PSNR/SSIM) 17.41/0.7769 21.88/0.8194 21.25/0.8453
-w PPU(PSNR/SSIM) 19.64/0.8157 21.23/0.8331 21.89/0.8570
-5. Conclusion
+Consequently, how to generate more accurate physical prior knowledge remains a challenge for future research.
+Regarding the extreme low-light conditions, we add a new set of qualitative results on the LSUI dataset.
+As shown in Figure 13, while our method is able to achieve noticeable overall enhancement in terms of brightness and color consistency, the recovery of fine-grained details remains challenging due to severe information loss.
+Moreover, the performance under such extreme conditions is further constrained by the limited quality of the available data, as the ground-truth images themselves are often noisy or poorly illuminated.
+Thus, we also plan to explore self-supervised learning paradigms and the construction of higher-quality datasets to better address extreme low-light underwater scenarios.
+V Conclusion
 In this paper, we develop a novel UIE framework, namely PA-Diff.
 With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
-To the best of our knowledge, it is the first the diffusion model based on physical perception in image enhancement tasks.
-Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into any deep learning network.
-PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.
+To the best of our knowledge, it is the first diffusion model based on physical perception for UIE tasks.
+Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into various deep learning network.
+PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments validate the effectiveness of each proposed component.